Above which number does 10% of z-scores fall?

Between what two z-scores does 95% of the data fall (go into the table- don't use the Empirical rule for this one.)?
Answer #1

solution:

a


P(Z > z ) =1 - 10%= 1 -0.10 =0.9

z= 1.28 using z table

b

middle 95% of score is

P(-z < Z < z) = 0.95

P(Z < z) - P(Z < -z) = 0.95

2 P(Z < z) - 1 = 0.95

2 P(Z < z) = 1 + 0. 95= 1.95

P(Z < z) = 1.95/ 2 = 0.975

P(Z <1.96 ) = 0.975

z  ±1.96